OpenStudy (anonymous):

What is the radius of the circle whose equation is x2 + y2 = 9? 3 9 81

OpenStudy (anonymous):

x2 + y2 = R^2 r^2 = 9 r = ?

OpenStudy (anonymous):

is it 3?

OpenStudy (whpalmer4):

Here's how you can check. If the radius is 3, then x=3 and y=0 should make a true number sentence. Does \[3^2+0^2=3^2\]?